OpenStudy (anonymous):

A coyote had 10 pups in her third litter. It was twice as many as in her second litter, and her second litter had three more pups than her first. Find the number of pups in her first litter. How would you put that in a equation form? please i need help!

OpenStudy (mertsj):

let x = # of pups in first litter Then x+3 = # of pups in second litter And 2(x+3) = pups in third litter which is given to be 10. So: 2(x+3) = 10 Divide both sides by 2: x+3 = 5 Subtract 3 from both sides: x = 2

OpenStudy (anonymous):

thanks, can you put it wall together for me in the form of a question

OpenStudy (mertsj):

Put what in the form of a question?

OpenStudy (mertsj):

Are you playing Jeopardy?

OpenStudy (anonymous):

no its mart of my math homework

OpenStudy (anonymous):

could you put the question in a form of a question. i have to read the question and put it in a equation and solve

OpenStudy (mertsj):

How many pups were in the first litter?

OpenStudy (anonymous):

the number of pups in the first litter. i know the answer i just need help to get it in the equation form.

OpenStudy (mertsj):

I gave you the equation form: 2(x+3)=10

OpenStudy (anonymous):

oh okay thankyou so much!